OpenStudy (anonymous):

Solvig rational equations (2c/c-4)-2=4/c+5

OpenStudy (anonymous):

Multiply through by common denominator (c - 4)(c + 5). Should turn it into a quadratic

jagr2713 (jagr2713):

\[(2\times \frac{ c }{ c }-4)-2=\frac{ 4 }{ c }+5\] ?? @ayyyyyyyyy25

OpenStudy (anonymous):

You need to eliminate the operations connecting to the variables by using inverse operations. what you do is put that on the other side of the equal sign.